In a Competitive examination of 5000 students, the marks of the examinees in statistics were found to be distributed normally with mean 45 and standard deviations 14.

Determine the number of examinees whose marks, out of 100 were;

(i) Less than 30.                                                         2MKS

(ii) Between 30 and 70.                                              2MKS

(iii) Between 60 and 80.                                          2MKS

(iv) More than 60.                                                 2MKS

(v) More than 40                              2MKS


1
Expert's answer
2022-03-11T05:53:51-0500

"\mu=45,\sigma=14,Z=(x-μ)/σ"

i. "P(x<30)=P(Z<(30-45)\/14)=P(Z<-1.07)"

From z distribution table

"P(Z<-1.07)=0.1423"

therefore number of students "=5000 * 0.1423 = 712"


ii. "P(30<x<70)=P((30-45)\/14<Z<(70-45)\/14)"

"=P(-1.07<Z<1.79)=P(Z<1.79)-P(Z<-1.07)"

"=0.9633-0.1423=0.821"

therefore number of students "=5000*0.821=4105"


iii. "P(60<x<80)=P((60-45)\/14<Z<(80-45)\/14)"

"=P(1.07<Z<2.5)=P(Z<2.5)-P(Z<1.07)"

"=0.9946-0.8577=0.1369"

therefore number of students "=5000*0.1369=685"


iv. "P(x>60)=P(Z>(60-45)\/14)=P(Z>1.07)=1-P(Z<1.07)"

"=1-0.8577=0.1423"

therefore number of students "=5000*0.1423=712"


v. "P(x>40)=P(Z>(40-45)\/14)=P(Z>-0.36)=1-P(Z<-0.36)"

"=1-0.3594=0.6406"

therefore number of students "=5000*0.6406=3203"